OpenStudy (anonymous):

WRITE THIS RATIO AS FRACTION IN LOWEST TERMS. $41.12 TO $31.20 (SIMPLIFY ANSWER . TYPE A FRACTION)

OpenStudy (anonymous):

\[\large \sf \frac{4112}{3120} \div \frac{2}{2}\]

OpenStudy (anonymous):

2056/1560

OpenStudy (anonymous):

\[\large \sf \frac{2056}{1560} \div \frac{2}{2}\]

OpenStudy (anonymous):

1028/780

OpenStudy (anonymous):

\[\large \sf \frac{1028}{780} \div \frac{2}{2}\]

OpenStudy (anonymous):

514/390

OpenStudy (anonymous):

\[\large \sf \frac{514}{390} \div \frac{2}{2}\]

OpenStudy (anonymous):

257/195

OpenStudy (anonymous):

why am i just dividing by 2 huh

OpenStudy (anonymous):

And this is as far as we can simplify

OpenStudy (anonymous):

Now let's convert them back to dollar form and we're done.

OpenStudy (anonymous):

ok lets do it

OpenStudy (anonymous):

Well it's pretty easy, just put a decimal in.

OpenStudy (anonymous):

but it has to be a fraction.. can fractions have decimals

OpenStudy (anonymous):

I'll do the top 257=$2.57

OpenStudy (anonymous):

$1.95

OpenStudy (anonymous):

Yup. So \[\large \sf \frac{$41.12}{$31.20}\] in its lowest term is \[\large \sf \frac{$2.57}{$1.95}\]